# Evolutionary Prediction Games

### v2

@@ D.5.1 Equivalent groups
Figure 8: Evolutionary dynamics of an oracle classifier in a three-group setting with equivalent groups (Section D.5.1). (Left) Data distributions of groups ${\left\{{A,B,C}\right\}}$, where $D_{B}=D_{A}$. Groups $A,B$ are oracle-equivalent, and group $C$ is not oracle-equivalent to $A$ or $B$ (Definition 9). (Right) Phase diagram of the corresponding replicator dynamics over the probability simplex (cf. Figure 6). The dynamics has three sets of fixed points, corresponding to Theorem 2: The convex hull of the singleton distributions corresponding to $A,B$ (stable), the singleton distribution of group $C$ (stable), and the convex hull of ${\bm{p}}\approx(0.45,0,0.55)$ and ${\bm{p}}^{\prime}\approx(0,0.45,0.55)$ (unstable).
Share the same optimal classifier and achieve the same optimal accuracy (formally satisfying Definition 9), then an extended version of Theorem 1 holds:
Theorem 2 (Oracle classifier with equivalent groups; Extension of Theorem 1).
@@
    Competitive exclusion: For all stable equilibria, $\operatorname*{support}({\bm{p}}^{*})$ is oracle-equivalent (Def. 9).
@@
In analogy to the proof of Theorem 1. The potential function remains $f({\bm{p}})=\mathrm{acc}_{p}(h_{\bm{p}})$, and so does the relation between extrema of the potential and equilibria of the dynamics. However, as groups are possibly equivalent, by Proposition 7 we obtain that $f({\bm{p}})$ may have additional weak maximizers in convex combinations of oracle-equivalent groups. ∎
@@ D.5.2 Heterogeneous fitness
Here we extend Theorem 1 to settings where retention functions $\nu_{k}$ vary across groups – See Theorem 3 below. For $K=2$, we show that a modified version of the theorem holds for any pair of continuous, strictly increasing retention functions. For $K>2$, a similar conclusion follows under positive affine retention functions $\nu_{k}(x)=ax-b_{k}$ with $a,b_{k}\geq 0$. In each case, we prove the extended theorem by constructing a transformed potential function and showing that it retains the properties required for convergence and equilibrium structure guarantees.
@@
Consider a two-group setting, and denote $[K]={\left\{{A,B}\right\}}$. Let $\mathcal{A}^{\mathrm{opt}}$ be an oracle classification algorithm. Denote ${\bm{p}}=(p,1-p)$, and denote $h_{p}\sim\mathcal{A}^{\mathrm{opt}}\left((p,1-p)\right)$.
Proposition 9.
@@
By Proposition 1, the average for loss for any fixed classifier is a convex combination of marginal loss functions. $h_{p}$ is optimal for $p$, and therefore:
@@
In contradiction to the optimality of $h_{p^{\prime}}$. Therefore, it must hold that $a\leq a^{\prime}$. Similarly, assume by contradiction that $b^{\prime}>b$ and obtain a similar contradiction. Figure 7 provides additional intuition. ∎
Proposition 10 (Potential function of a two-group population game; [43, Example 3.2.3]).
@@
Denote the transformed fitness function by $\tilde{F}_{A}((p,1-p))=\nu_{A}\left(\mathrm{acc}_{A}(h_{p})\right)$, and similarly denote $\tilde{F}_{B}((p,1-p))=\nu_{B}\left(\mathrm{acc}_{B}(h_{p})\right)$. Let $g(p)=\tilde{F}_{A}((p,1-p))-\tilde{F}_{B}((p,1-p))$, and denote by $\tilde{f}((p,1-p))$ the potential function for the transformed game, defined by Equation 8. By Proposition 9, it holds that $\tilde{F}_{A}((p,1-p))$ is increasing, $\tilde{F}_{B}((p,1-p))$ is decreasing, and therefore $g(p)$ is monotonically increasing as a function of $p$. By Equation 8, the potential $\tilde{f}$ is decreasing if $g(p)\leq 0$ for all $p$, increasing if $g(p)\geq 0$ for all $p$, and unimodal (decreasing and then increasing) if $g(p)$ changes signs. Note that $g(p)$ may change signs at most once due to monotonicity. In all of these cases, $g(p)$ doesn’t have a local maximizer at the interior of the simplex. ∎
@@
First, consider the game $F^{(a)}_{k}=aF_{k}({\bm{p}})$. By linearity of the gradient operation, the population game $F^{(a)}({\bm{p}})$ is a potential game with potential $f^{(a)}({\bm{p}})=af({\bm{p}})$. Then, consider the game $F^{(b)}_{k}({\bm{p}})=-b_{k}$. The game $F^{(b)}({\bm{p}})$ is a constant game, and therefore admits the linear potential function $f^{(b)}({\bm{p}})=-{\bm{b}}\cdot{\bm{p}}$ (see [43, Proposition 3.2.13]). Finally, by [43, Section 3.2.4], the game $\tilde{F}({\bm{p}})=F^{(a)}({\bm{p}})+F^{(b)}({\bm{p}})$ is a potential game with the potential function:
@@
By the proof of Theorem 1, $F({\bm{p}})$ is a potential game with potential function $f({\bm{p}})=\mathrm{acc}_{\bm{p}}(h_{\bm{p}})$. By Proposition 3 we obtain that $f({\bm{p}})$ is convex over the simplex. Multiplying a convex function by a positive constant and adding a linear function maintains convexity, and therefore $\tilde{f}({\bm{p}})$ is convex as well, and in particular does not have local maximizers in the interior of the simplex. ∎
Theorem 1 under heterogeneous retention.
In both cases analyzed above, we identify a modified potential function which does not have local minimizers in the interior of the simplex (Propositions 11 and 13). We now proceed to state a modified version of the central theorem:
Theorem 3 (Oracle classifier with heterogeneous retention; Extension of Theorem 1).
@@
In analogy to the proof of Theorem 1 (see Section D.5). The corresponding evolutionary prediction games are potential games, with Nash equilibria corresponding to local equilibria of the potential function, and stable equilibria corresponding to local maximizers of the potential function. By Propositions 11 and 13, potential functions for both families of retention functions don’t have local maximizers at the interior of the simplex, implying claims (2,3,4).
@@
and claim (1) follows from the fact that $\tilde{f}({\bm{p}})$ is a Lyapunov function. For monotone increasing functions, assume without loss of generality that $\tilde{F}_{A}({\bm{p}})\geq\tilde{F}_{B}({\bm{p}})$. Then from the positive correlation assumption $\dot{p}_{A}=-\dot{p}_{B}\geq 0$. From Proposition 9 we obtain that $\mathrm{acc}_{A}(h_{\bm{p}})$ is increasing and $\mathrm{acc}_{B}(h_{\bm{p}})$ is decreasing as a function of $p_{A}$. Combining the observations, we obtain:
@@ D.6 Soft-SVM
Definition 10 (Soft-SVM; [e.g., 45, Section 15.2]).
@@ D.6.1 Construction
Consider a two-group binary classification setting ($K=2,\mathcal{Y}={\left\{{-1,1}\right\}}$). Denote the two groups by $[K]={\left\{{A,B}\right\}}$. Denote by $\mathrm{Triangular}(a,b,c)$ the triangular distribution with parameters $(a,b,c)$ and probability density function:
@@
$Y|X$ =$=\begin{cases}1&X\geq 0\\ -1&X<0\end{cases}$
$D_{A}$ is the distribution of tuples $(X_{A},Y|X_{A})$, and $D_{B}$ is defined correspondingly as $D_{B}=(X_{B},Y|X_{B})$. The distributions are illustrated in Figure 3 (Left).
@@ D.6.2 Stable Mutualistic Coexistence
We show that the Soft-SVM algorithm can induce a stable, beneficial coexistence. We will consider learning in the population limit $n\to\infty$, where the loss minimization objective (Definition 10) converges towards its expected value:
Definition 11 (Soft-SVM in the population limit).
@@
Proof of Theorem 1.
Assume the Soft-SVM algorithm trains on a large dataset sampled from $D_{\bm{p}}$. Denote ${\bm{p}}=(1-p,p)$, and denote the probability density functions of groups ${\left\{{A,B}\right\}}$ by $f_{A}(x,y),f_{B}(x,y)$, respectively. Since the data is separable and labels are positively correlated with the feature $x$, it holds that $w^{*}>0$. Assume that regularization parameter $\gamma$ is large enough such that $w^{*}\leq 1$. From Proposition 14, it holds that:
@@
When $w\in(0,1]$, observe that the distribution defined in Section D.6.1 is suppported on $[-1,1]$, and therefore both integrals cover the whole distribution. We can therefore write:
@@
And as $\alpha>0.5$, it holds that $\left.\frac{\partial^{2}L(w,b)}{\partial p\partial b}\right|_{b=0,p=0.5}<0$. Let $\varepsilon>0$ which is sufficiently small. For $p^{\prime}=0.5+\varepsilon$ it holds that $\left.\frac{\partial L(w,b)}{\partial b}\right|_{b=0}<0$, and since $L(w,b)$ is convex, the optimal $b^{*}$ corresponding to $p^{\prime}$ satisfies $b^{*}>0$, and the decision boundary is smaller than 00. Assume that the deviation $\varepsilon$ is sufficiently small such that the new $b^{*}$ is also in the neighborhood of 00. By definition, in a sufficiently small neighborhood of zero, it holds by construction for each group:
@@ D.6.3 Bifurcations Induced by Regularization
In Figure 9, we vary the regularization strength for the distributions specified in Section D.6.1 with imbalance parameter $\alpha=0.75$. We observe that the system transition between three topologically-distinct phases: For weak regularization ($\lambda=10$), coexistence is beneficial but not stable. For intermediate regularization ($\lambda=100$), coexistence is beneficial and stable. Interestingly, for strong regularization ($\lambda=1000$), the game has five equilibria – Two stable single-group equilibria, one stable coexistence equilibia, and two unstable coexistence equilibria.
@@
Another practical approach to loss minimization is to use a class of models that fit the training data perfectly and then interpolate. Notable examples include the nearest-neighbor algorithm, and over-parameterized neural networks [e.g., 54]. Interpolation leads to perfect accuracy on the training set by definition, but often at the price of sensitivity to noise; as a concrete example, consider how 1-NN predictions change with the addition of a single point. The problem is that even in the limit, 1-NN cannot express the true $p(y|x)$, and remains sensitive to label noise. Our next result shows this can lead to coexistence:
Theorem 4.
@@
Formal proof in Section D.7.2. The construction is illustrated in Figure 3 (Center), and described formally in Section D.7.1. Informally, there are two groups defined symmetrically. Each group is composed of a majority and minority classes, and the majority class has label noise. For each group alone, the 1-NN algorithm makes perfect predictions on its minority class, and imperfect predictions on the majority class. Since groups are defined symmetrically, training on one group alone leads to better accuracy for the other—leading to stable coexistence. Moreover, Lemma 2 shows that for $k$-NN with $k\geq 3$, stable coexistence can also be beneficial for both groups.
Definition 12 (k-Nearest-Neighbors; e.g. [45, Section 19.1]).
@@ D.7.1 Construction
Consider a two-group binary classification setting ($K=2$, $\mathcal{Y}={\left\{{-1,1}\right\}}$). Denote the two groups by $[K]={\left\{{A,B}\right\}}$, and let $\alpha>0$, $\beta\geq 0.5$. The data distributions for the two groups are defined symmetrically:
@@
The construction is illustrated in Figure 3 (Center).
@@ D.7.2 Stable Coexistence
Proposition 15 (Convex combinations).
@@
Plugging eq. 12 into eq. 11, we obtain that the corresponding $q$ for the positive mixture is:
@@
Proposition 16 (Reflection-exchange symmetry).
@@
Proposition 17 (Expected prediction of $k$-NN).
@@
Lemma 2 (Expected accuracy).
@@
Where the identity $\mathbb{E}{\left[{y}\right]}=(1-2\alpha)$ is given by the definition of $\tilde{D}_{\mathrm{pos}}^{\alpha}$, the definition of $\gamma$ is given by Proposition 15, and the identity $\mathbb{E}{\left[{y_{i^{*}(x)}}\right]}=\left(-1+2\phi_{k}\left(\frac{\beta(1-p)}{\gamma}\alpha\right)\right)$ is given by Proposition 17.
@@
Figure 10: k-NN accuracy for the construction specified in Section D.7.1. The plots show alignment between the theoretical curves (Lemma 2) and empirical simulations (mean accuracies represented by dots). For $k=3,5$ the game has a stable beneficial coexistence.
Proposition 18.
@@
By applying Lemma 2. When $k=1$, the $\phi_{k}$ function satisfies:
@@
Plugging into eq. 13 and using the symmetry argument in Proposition 16, we obtain:
@@
Proof of Theorem 4.
By Proposition 18, using e.g. $(\alpha,\beta)=(0.2,0.8)$. We observe that the conditions of the proposition are satisfied for this choice of constants, as $\alpha=0.2<0.375=1-\frac{1}{2\cdot 0.8}=1-\frac{1}{2\beta}$. ∎

### v3

@@ D.5.1 Equivalent groups
Figure 9: Evolutionary dynamics of an oracle classifier in a three-group setting with equivalent groups (Section˜D.5.1). (Left) Data distributions of groups $\left\{{A,B,C}\right\}$}, where $D_{B}=D_{A}$. Groups $A,B$ are oracle-equivalent, and group $C$ is not oracle-equivalent to AA or $B$ (Definition˜D.9). (Right) Phase diagram of the corresponding replicator dynamics over the probability simplex (cf. Figure˜7). The dynamics has three sets of fixed points, corresponding to Theorem˜D.2: The convex hull of the singleton distributions corresponding to $A,B$ (stable), the singleton distribution of group $C$ (stable), and the convex hull of ${\bm{p}}\approx(0.45,0,0.55)$ and ${\bm{p}}^{\prime}\approx(0,0.45,0.55)$ (unstable).
Share the same optimal classifier and achieve the same optimal accuracy (formally satisfying Definition˜D.9), then an extended version of Theorem˜1 holds:
Theorem D.2 (Oracle classifier with equivalent groups; Extension of Theorem˜1).
@@
    Competitive exclusion: For all stable equilibria, $\operatorname*{support}({\bm{p}}^{*})$ is oracle-equivalent (Def. D.9).
@@
In analogy to the proof of Theorem˜1. The potential function remains $f({\bm{p}})=\mathrm{acc}_{p}(h_{\bm{p}})$, and so does the relation between extrema of the potential and equilibria of the dynamics. However, as groups are possibly equivalent, by Proposition˜D.7 we obtain that $f({\bm{p}})$ may have additional weak maximizers in convex combinations of oracle-equivalent groups. ∎
@@ D.5.2 Heterogeneous fitness
Here we extend Theorem˜1 to settings where retention functions $\nu_{k}$ vary across groups – See Theorem˜D.3 below. For $K=2$, we show that a modified version of the theorem holds for any pair of continuous, strictly increasing retention functions. For $K>2$, a similar conclusion follows under positive affine retention functions $\nu_{k}(x)=ax-b_{k}$ with $a,b_{k}\geq 0$. In each case, we prove the extended theorem by constructing a transformed potential function and showing that it retains the properties required for convergence and equilibrium structure guarantees.
@@
Consider a two-group setting, and denote $K]={\left\{{A,B}\right\}}$. Let $\mathcal{A}^{\mathrm{opt}}$ be an oracle classification algorithm. Denote ${\bm{p}}=(p,1-p)$, and denote $h_{p}\sim\mathcal{A}^{\mathrm{opt}}\left((p,1-p)\right)$.
Proposition D.9.
@@
By Proposition˜D.1, the average for loss for any fixed classifier is a convex combination of marginal loss functions. $h_{p}$ is optimal for $p$, and therefore:
@@
In contradiction to the optimality of $h_{p^{\prime}}$. Therefore, it must hold that $a\leq a^{\prime}$. Similarly, assume by contradiction that $b^{\prime}>b$ and obtain a similar contradiction. Figure˜8 provides additional intuition. ∎
Proposition D.10 (Potential function of a two-group population game; [52, Example 3.2.3]).
@@
Denote the transformed fitness function by $\tilde{F}_{A}((p,1-p))=\nu_{A}\left(\mathrm{acc}_{A}(h_{p})\right)$, and similarly denote $\tilde{F}_{B}((p,1-p))=\nu_{B}\left(\mathrm{acc}_{B}(h_{p})\right)$. Let $g(p)=\tilde{F}_{A}((p,1-p))-\tilde{F}_{B}((p,1-p))$, and denote by $\tilde{f}((p,1-p))$ the potential function for the transformed game, defined by Equation˜8. By Proposition˜D.9, it holds that $\tilde{F}_{A}((p,1-p))$ is increasing, $\tilde{F}_{B}((p,1-p))$ is decreasing, and therefore $g(p)$ is monotonically increasing as a function of $p$. By Equation˜8, the potential $\tilde{f}$ is decreasing if $g(p)\leq 0$ for all $p$, increasing if $g(p)\geq 0$ for all $p$, and unimodal (decreasing and then increasing) if $g(p)$ changes signs. Note that $g(p)$ may change signs at most once due to monotonicity. In all of these cases, $g(p)$ doesn’t have a local maximizer at the interior of the simplex. ∎
@@
First, consider the game $F^{(a)}_{k}=aF_{k}({\bm{p}})$. By linearity of the gradient operation, the population game $F^{(a)}({\bm{p}})$ is a potential game with potential $f^{(a)}({\bm{p}})=af({\bm{p}})$. Then, consider the game $F^{(b)}_{k}({\bm{p}})=-b_{k}$. The game $F^{(b)}({\bm{p}})$ is a constant game, and therefore admits the linear potential function $f^{(b)}({\bm{p}})=-{\bm{b}}\cdot{\bm{p}}$ (see [52, Proposition 3.2.13]). Finally, by [52, Section 3.2.4], the game $\tilde{F}({\bm{p}})=F^{(a)}({\bm{p}})+F^{(b)}({\bm{p}})$ is a potential game with the potential function:
@@
By the proof of Theorem˜1, $F({\bm{p}})$ is a potential game with potential function $f({\bm{p}})=\mathrm{acc}_{\bm{p}}(h_{\bm{p}})$. By Proposition˜D.3 we obtain that $f({\bm{p}})$ is convex over the simplex. Multiplying a convex function by a positive constant and adding a linear function maintains convexity, and therefore $\tilde{f}({\bm{p}})$ is convex as well, and in particular does not have local maximizers in the interior of the simplex. ∎
Theorem˜1 under heterogeneous retention.
In both cases analyzed above, we identify a modified potential function which does not have local minimizers in the interior of the simplex (Propositions D.11 and D.13). We now proceed to state a modified version of the central theorem:
Theorem D.3 (Oracle classifier with heterogeneous retention; Extension of Theorem˜1).
@@
In analogy to the proof of Theorem˜1 (see Section˜D.5). The corresponding evolutionary prediction games are potential games, with Nash equilibria corresponding to local equilibria of the potential function, and stable equilibria corresponding to local maximizers of the potential function. By Propositions D.11 and D.13, potential functions for both families of retention functions don’t have local maximizers at the interior of the simplex, implying claims (2,3,4).
@@
and claim (1) follows from the fact that $\tilde{f}({\bm{p}})$ is a Lyapunov function. For monotone increasing functions, assume without loss of generality that $\tilde{F}_{A}({\bm{p}})\geq\tilde{F}_{B}({\bm{p}})$. Then from the positive correlation assumption $\dot{p}_{A}=-\dot{p}_{B}\geq 0$. From Proposition˜D.9 we obtain that $\mathrm{acc}_{A}(h_{\bm{p}})$ is increasing and $\mathrm{acc}_{B}(h_{\bm{p}})$ is decreasing as a function of $p_{A}$. Combining the observations, we obtain:
@@ D.6 Soft-SVM
Definition D.10 (Soft-SVM; [e.g., 54, Section 15.2]).
@@ D.6.1 Construction
Consider a two-group binary classification setting ($K=2,\mathcal{Y}={\left\{{-1,1}\right\}}$). Denote the two groups by $K]={\left\{{A,B}\right\}}$. Denote by $\mathrm{Triangular}(a,b,c)$ the triangular distribution with parameters $(a,b,c)$ and probability density function:
@@
$Y|X$ $1X<0\displaystyle$=
$D_{A}$ is the distribution of tuples $(X_{A},Y|X_{A})$, and $D_{B}$ is defined correspondingly as $D_{B}=(X_{B},Y|X_{B})$. The distributions are illustrated in Figure˜3 (Left).
@@ D.6.2 Stable Mutualistic Coexistence
We show that the Soft-SVM algorithm can induce a stable, beneficial coexistence. We will consider learning in the population limit $n\to\infty$, where the loss minimization objective (Definition˜D.10) converges towards its expected value:
Definition D.11 (Soft-SVM in the population limit).
@@
Proof of Theorem˜2.
Assume the Soft-SVM algorithm trains on a large dataset sampled from $D_{\bm{p}}$. Denote ${\bm{p}}=(1-p,p)$, and denote the probability density functions of groups $\left\{{A,B}\right\}$} by $f_{A}(x,y),f_{B}(x,y)$, respectively. Since the data is separable and labels are positively correlated with the feature $x$, it holds that $w^{*}>0$. Assume that regularization parameter $\gamma$ is large enough such that $w^{*}\leq 1$. From Proposition˜D.14, it holds that:
@@
When $w\in(0,1]$, observe that the distribution defined in Section˜D.6.1 is suppported on [1,1][-1,1], and therefore both integrals cover the whole distribution. We can therefore write:
@@
And as $\alpha>0.5$, it holds that $\left.\frac{\partial^{2}L(w,b)}{\partial p\partial b}\right|_{b=0,p=0.5}<0$. Let $\varepsilon>0$ which is sufficiently small. For $p^{\prime}=0.5+\varepsilon$ it holds that $\left.\frac{\partial L(w,b)}{\partial b}\right|_{b=0}<0$, and since $L(w,b)$ is convex, the optimal $b^{*}$ corresponding to $p^{\prime}$ satisfies $b^{*}>0$, and the decision boundary is smaller than 0. Assume that the deviation $\varepsilon$ is sufficiently small such that the new $b^{*}$ is also in the neighborhood of 0. By definition, in a sufficiently small neighborhood of zero, it holds by construction for each group:
@@ D.6.3 Bifurcations Induced by Regularization
In Figure˜10, we vary the regularization strength for the distributions specified in Section˜D.6.1 with imbalance parameter $\alpha=0.75$. We observe that the system transition between three topologically-distinct phases: For weak regularization ($\lambda=10$), coexistence is beneficial but not stable. For intermediate regularization ($\lambda=100$), coexistence is beneficial and stable. Interestingly, for strong regularization ($\lambda=1000$), the game has five equilibria – Two stable single-group equilibria, one stable coexistence equilibia, and two unstable coexistence equilibria.
@@
Another practical approach to loss minimization is to use a class of models that fit the training data perfectly and then interpolate. Notable examples include the nearest-neighbor algorithm, and over-parameterized neural networks [e.g., 64]. Interpolation leads to perfect accuracy on the training set by definition, but often at the price of sensitivity to noise; as a concrete example, consider how 1-NN predictions change with the addition of a single point. The problem is that even in the limit, 1-NN cannot express the true $p(y|x)$, and remains sensitive to label noise. Our next result shows this can lead to coexistence:
Theorem D.4.
@@
Formal proof in Section˜D.7.2. The construction is illustrated in Figure˜3 (Center), and described formally in Section˜D.7.1. Informally, there are two groups defined symmetrically. Each group is composed of a majority and minority classes, and the majority class has label noise. For each group alone, the 1-NN algorithm makes perfect predictions on its minority class, and imperfect predictions on the majority class. Since groups are defined symmetrically, training on one group alone leads to better accuracy for the other—leading to stable coexistence. Moreover, Lemma˜D.2 shows that for $k$-NN with $k\geq 3$, stable coexistence can also be beneficial for both groups.
Definition D.12 (k-Nearest-Neighbors; e.g. [54, Section 19.1]).
@@ D.7.1 Construction
Consider a two-group binary classification setting ($K=2$, $\mathcal{Y}={\left\{{-1,1}\right\}}$). Denote the two groups by $K]={\left\{{A,B}\right\}}$, and let $\alpha>0$, $\beta\geq 0.5$. The data distributions for the two groups are defined symmetrically:
@@
The construction is illustrated in Figure˜3 (Center).
@@ D.7.2 Stable Coexistence
Proposition D.15 (Convex combinations).
@@
Plugging eq.˜12 into eq.˜11, we obtain that the corresponding $q$ for the positive mixture is:
@@
Proposition D.16 (Reflection-exchange symmetry).
@@
Proposition D.17 (Expected prediction of $k$-NN).
@@
Lemma D.2 (Expected accuracy).
@@
Where the identity $\mathbb{E}{\left[{y}\right]}=(1-2\alpha)$ is given by the definition of $\tilde{D}_{\mathrm{pos}}^{\alpha}$, the definition of $\gamma$ is given by Proposition˜D.15, and the identity $\mathbb{E}{\left[{y_{i^{*}(x)}}\right]}=\left(-1+2\phi_{k}\left(\frac{\beta(1-p)}{\gamma}\alpha\right)\right)$ is given by Proposition˜D.17.
@@
Figure 11: k-NN accuracy for the construction specified in Section˜D.7.1. The plots show alignment between the theoretical curves (Lemma˜D.2) and empirical simulations (mean accuracies represented by dots). For $k=3,5$ the game has a stable beneficial coexistence.
Proposition D.18.
@@
By applying Lemma˜D.2. When $k=1$, the $\phi_{k}$ function satisfies:
@@
Plugging into eq.˜13 and using the symmetry argument in Proposition˜D.16, we obtain:
@@
Proof of Theorem˜D.4.
By Proposition˜D.18, using e.g. $(\alpha,\beta)=(0.2,0.8)$. We observe that the conditions of the proposition are satisfied for this choice of constants, as $\alpha=0.2<0.375=1-\frac{1}{2\cdot 0.8}=1-\frac{1}{2\beta}$. ∎